OpenStudy (anonymous):

The sum of two consecutive integers is at least 36. What is the least possible pair of integers? A. 17 and 18 B. 19 and 20 C. 18 and 19 D. 16 and 17

OpenStudy (anonymous):

"Sum" means 'add'. 'Consecutive' means they are in order without skipping one. 2,3 are consecutive numbers. 2,4 are not consecutive. Add each set of two numbers to see which ones are 36 or higher. Of all the ones that do add to 36 or more, they want the smallest pair of numbers as the answer.

OpenStudy (anonymous):

so then would it be c?

OpenStudy (anonymous):

A. 17 + 18 = 35 B. 19 + 20 = 39 C. 18 + 19 = 37 D. 16 + 17 = 33 Only two sets are equal to 36 or more. Yes C is the answer.

OpenStudy (anonymous):

Yes. C. is correct.

OpenStudy (anonymous):

ok. thank you for your help! :)

OpenStudy (anonymous):

Your welcome!

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

i have another question

OpenStudy (anonymous):

\[5\le \frac{ x }{ -2 }\]

OpenStudy (anonymous):

\[-7\ge13-5x\] this is my next one too. so I need help with both.

OpenStudy (anonymous):

both of the last two questions i put i need help with

OpenStudy (anonymous):

nvm i got them. you took too long. jk! ;P

OpenStudy (anonymous):

Sorry. For 5 <= x/(-2) Did you remember to switch the direction of the inequality sign after you multiplied both sides by (-2)?

OpenStudy (anonymous):

Answer would be x is less than or equal to -10.

OpenStudy (anonymous):

For the next one: -7>=13-5x Add 5x to both sides to get: -7-5x>=13 Then add 7 to both sides to get: -5x>=13+7 So -5x>=20 Next you divide both sides by -5 BUT you have to remember to switch the direction of the inequality sign! So x<=-4

OpenStudy (anonymous):

ok thank you! I got it before you typed the answer! xD you took to long! lol! jk

OpenStudy (anonymous):

Sorry I left the room to go do some stuff. I just check every now and then.

OpenStudy (anonymous):

oh ok. well thanks for your help. that is all from me for now

OpenStudy (anonymous):

Ok. Bye.

OpenStudy (anonymous):

see ya